OpenStudy (anonymous):

the slope between the two points (-3, 1) and (3, -6

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

the slope formula is as follows: slope = (y2-y1)/(x2-x1) you have two points (x1,y1) in (-3,1) (x2,y2) in (3,-6) plug it in and wallah you have the slope

OpenStudy (anonymous):

-3-3 and 1-3?

OpenStudy (anonymous):

-6-3?

hartnn (hartnn):

not actually, lets take it one by one, y1-y2 = 1-(-6)=...?

OpenStudy (anonymous):

7

hartnn (hartnn):

yes, that was numerator, now denominator =x1-x2 = -3-3 =... ?

OpenStudy (anonymous):

-6

hartnn (hartnn):

correct, so your slope will be ...?

OpenStudy (anonymous):

-7/6

hartnn (hartnn):

you are \(\huge \color{red}{\checkmark}\)

OpenStudy (anonymous):

sweet thanks

hartnn (hartnn):

welcome ^_^